1) Sue and Bob recently purchased a brick house for $150,000. They financed that amount at 8
percent for 30 years. The tax rate in their area is 71.57 mills and the rate of assessment is 40
percent. They purchased a homeowners policy for the purchase price of the house. The fire
protection in the neighborhood is rated class 9. For the month of July, they recorded the
following: $69.20 for electricity, $44.85 for telephone, $39.95 for cable TV, $18.80 for water,
and $74.65 to repair a door. They have a combined monthly net income of $5,400. What is their
monthly mortgage payment? What is the monthly insurance premium? What are their monthly
taxes? What was their total monthly housing cost for July? Is it within the FHA
recommendation?
1
Expert's answer
2020-04-22T14:16:57-0400

1.Monthly mortgage payment:

"X=S\times\frac{\frac{i}{12}}{1-(1+\frac{i}{12})^-{n\times12}}=150 000\times\frac{0.007}{1-(1+0.007)^-360}=1 155.12"


since the rate of 8 percent is annual, then divide this annual rate by 12

"i=\frac{i}{12}=\frac{0.08}{12}=0.007"

since the number is 30 years old, then the number of month

"30\times12=360"

2.Monthly insurance premium:

"150 000\times0.09=13 500"

since the fire protection class is 9, the percentage for calculating insurance will be 9%, 0.09

"\frac{13500}{360}=37.5"

3.Monthly tax:

"150 000\times0.4=60 000"

calculation of property value

"\frac{60 000}{360}=166.67"

4.Total cost of housing

utilities July: 69.20+44.85+39.95+18.80+74.65=247.45

1 155.12+37.5+166.67+247.45=1606.74

5400-1606.74=3793.26

"\frac{1606.74}{5400}=0.2975" , 29,75% is debt-to-income ratio

total cost is 29.75% of net income per month, the couple has the opportunity to cope with a mortgage.The FHA recommendation is debt-to-income ratio of 50% or less.

And in this case, the ratio of 29.75%